# Relative periodic solutions in spatial Kepler problem with symmetric perturbation

### v1

@@ 2. The Kepler Problem
In both cases, $p_{z}(t)<0$ for $t\in(0,t_{2})$, and the plane formed by major axis direction of the ellipse orbit and angular momentum direction contains $z$-axis. When $\xi(t)$ intersects the $r$-axis (i.e., at $t=t_{1}$), $\hat{\theta}(t_{1})=\pi/2$. Thus, every trajectory starting from Hill’s boundary is a brake orbit.
For a solution $\bar{\xi}(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ with $\bar{\xi}(0)\in\Sigma_{+}\cap\{p_{r}=0\}$, Lemma 2.2 ensures a minimal $t_{1}>0$ such that $\bar{\xi}(t_{1})\in\Sigma_{-}$. Since $\bar{\xi}(t)$ is an elliptic orbit, symmetry implies its major axis on the $(x,y)$-plane and $\bar{\xi}(t_{1})\in\Sigma_{-}\cap\{p_{r}=0\}$. Consequently, $\bar{\xi}(t)$ is a $z$-symmetric periodic orbit.
@@
For a brake orbit $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$,
@@
Interpreting $\xi(t)$ as a solution to the planar Kepler problem,
@@
From (27) and $p_{\varphi}=\rho^{2}\dot{\varphi}=0$, we derive
@@
To determine the sign of $\tilde{p}_{r}$, observe that near pericenter, $\tilde{\rho}(t)$ increases. Thus, $\tilde{p}_{r}=\tilde{p}_{\rho}>0$, giving
@@
Assume $\xi(t)$ is a brake orbit. It also a $z$-symmetric orbit only when $p_{r}(t_{1})=0$. By lemma 2.3, we get the unique root $r(0)=\frac{\varpi}{\sqrt{-2h}}$. So there exists one and only one $z$-symmetric brake orbit $\xi_{bz}(t)$ besides $\xi_{0}(t)$ within $\mathcal{M}$. Due to these two orbits linked together exactly once, they formed a hopf link. ∎
@@
The orbit $\xi_{bz}(t)$ corresponds to a circular solution of (9) with angular momentum $\varpi$ about $z$-axis.
@@ 3. The Kepler Problem with a symmetric perturbation
By Proposition 2.1, when $2h\varpi^{2}\in(-1,0)$, $\mathcal{M}(h,\varpi,0)$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$, where
@@
whose solutions is $\xi_{\varepsilon}(t)$. ∎
@@
We say a Reeb orbit $\xi$ is elliptic if all eigenvalues of its monodromy matrix $\gamma$ lie on $\mathbb{U}$ and semi-simple, where $\mathbb{U}$ denotes the unit circle in the complex plane. If, additionally, the angle $\vartheta$ is irrational for eigenvalues $\sigma=e^{i\vartheta}$ and $\sigma^{-1}=e^{-i\vartheta}$, we say $\xi$ is irrationally elliptic. A Reeb orbit $\xi$ is hyperbolic if all eigenvalues of its monodromy matrix $\gamma$ are real and distinct from $\pm 1$.
@@ 3.1. Estimate the contact volume and action value
When $(\varpi,h)\in\Gamma_{\varepsilon}$, let $r(\theta)=\varpi^{2}/(1+e\cos\theta)$, $e=\sqrt{1+2h\varpi^{2}}$, we have
@@
Let $F(x,y,\varepsilon)\in C^{\infty}(\mathbb{R}^{2}\times(-\varepsilon_{0},\varepsilon_{0}),\mathbb{R})$ and $D\subset\mathbb{R}^{2}$ is a closed region, $\partial D$ is the boundary of $D$. Assume $\phi^{\varepsilon}$ is a family of differential homeomorphism, $F(x,y,0)|_{\partial D}=0$, and $\left.\frac{\partial F}{\partial x}(x,y,0)\right|_{\partial D}\neq 0$, $\left.\frac{\partial F}{\partial x}(x,y,0)\right|_{\partial D}\neq 0$. then
@@
Here $\mathcal{H}_{\varepsilon}$ is the Hill’s region of $\mathcal{M}_{\varepsilon}$ and $\mathcal{H}$ is the Hill’s region of $\mathcal{M}$.
@@
Here we set $r_{1,\varepsilon},r_{2,\varepsilon}$ are the roots of equation $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h$. ∎
@@ 3.2. Seifert rotation number of $\xi_{\varepsilon}$
The definitions and properties of the Maslov-type index (also known as the Conley-Zehnder index) and the mean index are listed in Appendix, for more details, we refer the readers to [15]. We denote $\text{Rot}(\xi_{\varepsilon})$ by $\text{Rot}_{\varepsilon}$ below.
@@
Using $\theta$ as the time variable, let $T_{\varepsilon}$ denote the period of $\xi_{\varepsilon}(\theta)$. Recall that $r(\theta)=\varpi^{2}/(1+e\cos\theta)$, and let $r_{1,\varepsilon}$ and $r_{2,\varepsilon}$ be the roots of $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h.$ From $r^{2}\dot{\theta}=\varpi$, the period $T_{\varepsilon}$ is expressed as
@@
(46) $\nabla K_{\varpi,\varepsilon}=\frac{\rho^{2}}{\varpi}\nabla H_{\varpi,\varepsilon}+\nabla(\rho^{2}/\varpi)(H_{\varpi,\varepsilon}-h),$
@@
For $\xi_{\varepsilon}(\tau)$ restricted to the $\varphi$-axis ($\varphi(\xi_{\varepsilon})=0$)
@@
For a $2\times 2$ symplectic path from mechanical systems, the analysis is simpler than the general case. Let $\sigma_{\varepsilon}(\tau)=\rho(\tau)e^{i\vartheta(\tau)}$ and $\sigma_{\varepsilon}^{-1}(\tau)$ denote the eigenvalues of $\gamma_{\varepsilon}(\tau)$, where $\vartheta(t)$ is a continuous non-decreasing function. The mean index is then given by $\hat{i}(\gamma_{\varepsilon}(\tau))=\frac{\vartheta(T_{\varepsilon})-\vartheta(0)}{\pi}.$ However, directly calculating the derivative of rotation number $\text{Rot}_{\varepsilon}$ remains challenging. To address this, we utilize the trace formula introduced by the first author and collaborators [12]. This framework was originally introduced to analyze the linear stability of the Lagrange elliptic relative equilibrium. Below, we briefly summarize relevant results for our purposes. Let $T>0$, $I_{2n}$ denote the $2n\times 2n$ identity matrix, and $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$ be the standard symplectic matrix. Let $\mathcal{B}(k)$ denote the space of $k\times k$ symmetric matrices, and let $B(t),D(t)\in C([0,T],\mathcal{B}(2n))$. For the anti-periodic linear Hamiltonian system
@@
Define $D(f)$ and $E(f)$ as below
@@
Differentiating twice with respect to $\varepsilon$ at $\varepsilon=0$,
@@
(52) $D(f)/\varpi^{4}=(\sigma_{0}^{\prime})^{2}.$
@@
(53) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\frac{E(f)}{2\pi\varpi^{2}}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
@@
(54) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
Repeating the eigenvalue analysis for (54), the monodromy matrix eigenvalues are real. By Sturm’s comparison theorem: For $E(f)>0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}>0$. For $E(f)<0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}<0$. ∎
@@
Although D(f)=D(f)D(f)=D(-f), the derivative of $\text{Rot}_{\varepsilon}$ is opposite.
@@
$\xi_{\varepsilon}(t)$ is elliptic, when $D(f)>0$, and hyperbolic, when $D(f)<0$.
@@
We say $\xi_{\varepsilon}(t)$ is linear stable, if $\xi_{\varepsilon}(t)$ is elliptic, and linear unstable, if $\xi_{\varepsilon}(t)$ is hyperbolic.
@@
    $D(f)\geq 0$ and $\tilde{V}(f)\neq 2C(e)\tilde{A}(f)+C^{2}(e)\text{sign}(E(f))\sqrt{D(f)}/2$.
@@
$\left.\frac{\mathrm{d}\text{Vol}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\neq 2A(\xi_{r}(t))\left.\frac{\mathrm{d}A_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}-A^{2}(\xi_{r}(t))\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0},$
then infinitely many periodic orbits exist for sufficiently small $\varepsilon$. By Theorem 3.3, Lemma 3.3, and equation (19), this inequality is equivalent to conditions (ii). Furthermore, by Corollary 3.1, Theorem 3.1, and Theorem 3.2, condition (i) implies that $\xi_{\varepsilon}$ is hyperbolic and that infinitely many periodic orbits exist. ∎
@@
From previous definition, $\tilde{V}(f)$, $\tilde{A}(f)$ and $\sqrt{D(f)}$ are the scaled derivative of $\text{Vol}_{\varepsilon}$, $A_{\varepsilon}$ and $\text{Rot}_{\varepsilon}$ correspondingly.
While one can obtain infinitely many periodic orbits with the above theorem, it requires some additional conditions besides the perturbation is small enough. Meanwhile our next result shows the existence of at least one periodic orbit when the symmetric perturbation is small enough without any further assumption.
@@
By the IFT, there exists a unique $\tilde{T}(r,\varepsilon)$ near $(r,0,T(r))$ such that $\tilde{T}(r,0)=T(r)$ and
@@
Applying the IFT again, there exists a unique $\tilde{r}(\varepsilon)$ near $(0,r_{1})$ such that $\tilde{r}(0)=r_{1}$ and
@@
There may be many even infinitely many $z$-symmetric brake orbits, but only $\xi_{bz,\varepsilon}(t)$ forms a Hopf link with $\xi_{\varepsilon}(t)$.

### v2

@@ 2. The Kepler Problem
In both cases, $p_{z}(t)<0$ for $t\in(0,t_{2})$, and the plane formed by the major axis direction of the elliptic orbit and angular momentum direction contains $z$-axis. When $\xi(t)$ intersects the $r$-axis (i.e., at $t=t_{1}$), $\hat{\theta}(t_{1})=\pi/2$. Thus, every trajectory starting from Hill’s boundaries is a brake orbit.
For a solution $\bar{\xi}(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ with $\bar{\xi}(0)\in\Sigma_{+}\cap\{p_{r}=0\}$, Lemma 2.2 ensures a minimal $t_{1}>0$ such that $\bar{\xi}(t_{1})\in\Sigma_{-}$. Since $\bar{\xi}(t)$ is an elliptic orbit, symmetry implies that its major axis lies on the $(x,y)$-plane and $\bar{\xi}(t_{1})\in\Sigma_{-}\cap\{p_{r}=0\}$. Consequently, $\bar{\xi}(t)$ is a $z$-symmetric periodic orbit.
@@
For a brake orbit $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ with $\Pi_{r,z}\xi(0)\in\mathcal{B}_{+}$, we have
@@
Interpreting $\xi(t)$ as a solution of the planar Kepler problem,
@@
From (27) and $p_{\varphi}=\rho^{2}\dot{\varphi}=0$, we obtain
@@
To determine the sign of $\tilde{p}_{r}$, observe that near the pericenter, $\tilde{\rho}(t)$ increases. Thus, $\tilde{p}_{r}=\tilde{p}_{\rho}>0$, giving
@@
Assume $\xi(t)$ is a brake orbit. It is also a $z$-symmetric orbit if and only if $p_{r}(t_{1})=0$. By Lemma 2.3, we get the unique root $r(0)=\frac{\varpi}{\sqrt{-2h}}$. So there exists one and only one $z$-symmetric brake orbit $\xi_{bz}(t)$ besides $\xi_{0}(t)$ within $\mathcal{M}$. Due to these two orbits linked together exactly once, they form a Hopf link. ∎
@@
The orbit $\xi_{bz}(t)$ corresponds to a circular solution of (9) with angular momentum $\varpi$ about the $z$-axis.
@@ 3. The Kepler Problem with a symmetric perturbation
By Proposition 2.1, when $2h\varpi^{2}\in(-1,0)$, $\mathcal{M}(h,\varpi,0)$ is a compact manifold diffeomorphic to $\mathbb{S}^{3}$, where
@@
whose solution is $\xi_{\varepsilon}(t)$. ∎
@@
We say a Reeb orbit $\xi$ is elliptic if all eigenvalues of its monodromy matrix $\gamma$ lie on $\mathbb{U}$ and are semi-simple, where $\mathbb{U}$ denotes the unit circle in the complex plane. If, additionally, the angle $\vartheta$ is irrational for the eigenvalues $\sigma=e^{i\vartheta}$ and $\sigma^{-1}=e^{-i\vartheta}$, we say $\xi$ is irrationally elliptic. A Reeb orbit $\xi$ is hyperbolic if all eigenvalues of its monodromy matrix $\gamma$ are real and distinct from $\pm 1$.
@@ 3.1. Estimate the contact volume and action value
When $(\varpi,h)\in\Gamma_{\varepsilon}$, let $r(\theta)=\varpi^{2}/\penalty 50(1+e\cos\theta)$, $e=\sqrt{1+2h\varpi^{2}}$, we have
@@
Let $F(x,y,\varepsilon)\in C^{\infty}(\mathbb{R}^{2}\times(-\varepsilon_{0},\varepsilon_{0}),\mathbb{R})$ and $D\subset\mathbb{R}^{2}$ be a closed domain, $\partial D$ is the boundary of $D$. Assume $\phi^{\varepsilon}$ is a family of diffeomorphisms, $F(x,y,0)|_{\partial D}=0$, and $\left.\frac{\partial F}{\partial x}(x,y,0)\right|_{\partial D}\neq 0$, $\left.\frac{\partial F}{\partial y}(x,y,0)\right|_{\partial D}\neq 0$, then
@@
Here $\mathcal{H}_{\varepsilon}$ is Hill’s region of $\mathcal{M}_{\varepsilon}$ and $\mathcal{H}$ is Hill’s region of $\mathcal{M}$.
@@
Here we set $r_{1,\varepsilon}$ and $r_{2,\varepsilon}$ are the roots of the equation $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h$. ∎
@@ 3.2. Seifert rotation number of $\xi_{\varepsilon}$
The definitions and properties of the Maslov-type index (also known as the Conley-Zehnder index) and the mean index are listed in Appendix, for more details, we refer the reader to [15]. We denote $\text{Rot}(\xi_{\varepsilon})$ by $\text{Rot}_{\varepsilon}$ below.
@@
Using $\theta$ as the time variable, let $T_{\varepsilon}$ denote the period of $\xi_{\varepsilon}(\theta)$. Recall that $r(\theta)=\varpi^{2}/\penalty 50(1+e\cos\theta)$, and let $r_{1,\varepsilon}$ and $r_{2,\varepsilon}$ be the roots of $\frac{\varpi^{2}}{2r^{2}}-\frac{1}{r}+\varepsilon f(r,0,\varepsilon)=h.$ From $r^{2}\dot{\theta}=\varpi$, the period $T_{\varepsilon}$ is expressed as
@@
(46) $\nabla K_{\varpi,\varepsilon}=\frac{\rho^{2}}{\varpi}\nabla H_{\varpi,\varepsilon}+\nabla(\rho^{2}/\penalty 50\varpi)(H_{\varpi,\varepsilon}-h),$
@@
For $\xi_{\varepsilon}(\tau)$ restricted to the $\varphi$-axis ($\varphi(\xi_{\varepsilon})=0$), we have
@@
For a $2\times 2$ symplectic path from mechanical systems, the analysis is simpler than in the general case. Let $\sigma_{\varepsilon}(\tau)=\rho(\tau)e^{i\vartheta(\tau)}$ and $\sigma_{\varepsilon}^{-1}(\tau)$ denote the eigenvalues of $\gamma_{\varepsilon}(\tau)$, where $\vartheta(t)$ is a continuous non-decreasing function. The mean index is then given by $\hat{i}(\gamma_{\varepsilon}(\tau))=\frac{\vartheta(T_{\varepsilon})-\vartheta(0)}{\pi}.$ However, directly calculating the derivative of rotation number $\text{Rot}_{\varepsilon}$ remains challenging. To address this, we utilize the trace formula introduced by the first author and collaborators [12]. This framework was originally developed to analyze the linear stability of the Lagrange elliptic relative equilibria. Below, we briefly summarize relevant results for our purposes. Let $T>0$, $I_{2n}$ denote the $2n\times 2n$ identity matrix, and $J=\begin{pmatrix}0&-I_{n}\\ I_{n}&0\end{pmatrix}$ be the standard symplectic matrix. Let $\mathcal{B}(k)$ denote the space of $k\times k$ symmetric matrices, and let $B(t),D(t)\in C([0,T],\mathcal{B}(2n))$. For the anti-periodic linear Hamiltonian system
@@
Define $D(f)$ and $E(f)$ as below, where $\tilde{T}(f)$ is introduced in (43),
@@
Differentiating twice with respect to $\varepsilon$ at $\varepsilon=0$, we have
@@
(52) $D(f)/\penalty 50\varpi^{4}=(\sigma_{0}^{\prime})^{2}.$
@@
(53) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\frac{E(f)}{2\pi\varpi^{2}}I_{2}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}I_{2}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
@@
(54) $\left\{\begin{aligned} &\tilde{\eta}_{2}^{\prime}=J\left(I_{2}+\varepsilon\left(D_{\varepsilon}-\frac{E(f)}{2\pi\varpi^{2}}I_{2}\right)\right)\tilde{\eta}_{2},\\ &\tilde{\eta}_{2}(0)=-\tilde{\eta}_{2}(2\pi).\end{aligned}\right$.
Repeating the eigenvalue analysis for (54), the monodromy matrix eigenvalues are real. By Sturm’s Comparison Theorem: for $E(f)>0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}>0$; for $E(f)<0$, $\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}<0$. ∎
@@
Although D(f)=D(f)D(f)=D(-f), the derivative of $\text{Rot}_{\varepsilon}$ is opposite in sign.
@@
$\xi_{\varepsilon}(t)$ is elliptic when $D(f)>0$ and hyperbolic when $D(f)<0$.
@@
We say $\xi_{\varepsilon}(t)$ is linearly stable, if $\xi_{\varepsilon}(t)$ is elliptic, and linearly unstable, if $\xi_{\varepsilon}(t)$ is hyperbolic.
@@
[…]
where $D(f)$, $E(f)$ are defined in Theorem 3.3, and $\tilde{V}(f)$, $\tilde{A}(f)$ are defined in Lemma 3.3.
@@
$\left.\frac{\mathrm{d}\text{Vol}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}\neq 2A(\xi_{0}(t))\left.\frac{\mathrm{d}A_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0}-A^{2}(\xi_{0}(t))\left.\frac{\mathrm{d}\text{Rot}_{\varepsilon}}{\mathrm{d}\varepsilon}\right|_{\varepsilon=0},$
then infinitely many periodic orbits exist for sufficiently small $\varepsilon$. By Theorem 3.3, Lemma 3.3, and equation (19), this inequality is equivalent to condition (ii). Furthermore, by Corollary 3.1, Theorem 3.1, and Theorem 3.2, condition (i) implies that $\xi_{\varepsilon}$ is hyperbolic and that infinitely many periodic orbits exist. ∎
@@
From previous definition, $\tilde{V}(f)$, $\tilde{A}(f)$ and $\sqrt{D(f)}$ are the scaled derivatives of $\text{Vol}_{\varepsilon}$, $A_{\varepsilon}$ and $\text{Rot}_{\varepsilon}$ respectively.
While one can obtain infinitely many periodic orbits using the above theorem, it requires some additional conditions besides the perturbation being small enough. Meanwhile, our next result shows the existence of at least one periodic orbit when the symmetric perturbation is small enough without any further assumption.
@@
By the IFT, there exists a unique $\tilde{T}(r,\varepsilon)$ near $(T(r),r,0)$ such that $\tilde{T}(r,0)=T(r)$ and
@@
Applying the IFT again, there exists a unique $\tilde{r}(\varepsilon)$ near $(r_{1},0)$ such that $\tilde{r}(0)=r_{1}$ and
@@
There may even be infinitely many $z$-symmetric brake orbits, but only $\xi_{bz,\varepsilon}(t)$ forms a Hopf link with $\xi_{\varepsilon}(t)$.